what is slope of line (1,-1) (4,-1) how do you find this answer
Answer by vksarvepalli(154)

the slope(m) of a line passing through two points (x1,y1) and (x2,y2) is given by
slope m = (y2-y1)/(x2-x1)
so, slope of the line through the two points (1,-1) and (4,-1) is
slope = (-1+1)/(4-1) = (0)/(3)
so slope = 0 answer